# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 3.3 Nearly-Linear Time Condensers
The downside of Theorem 3.10 is that it requires the entropy in the source to be $\Omega(\log^{2}(1/\varepsilon))$, instead of the optimal $\Omega(\log(1/\varepsilon))$. Instead, we can use a lossy condenser [^15] Our extractor will lose a small constant fraction of the entropy, so losing a small constant fraction of the entropy in the condensing step will not make much difference. based on Reed–Solomon codes. Unfortunately, this comes at the expense of computing a generator of a field of size $\operatorname{poly}(1/\varepsilon)$, which we do not know how to do in nearly-linear time for arbitrary $\varepsilon$-s. We consider it a one-time preprocessing step, as it does not depend on the inputs to the condenser.
Theorem 3.11 (the lossy RS condenser, [GUV09]).
@@
For any constant $\alpha\in(0,1)$ the following holds, for every $n\in\mathbb{N}$, and any $0<\varepsilon\leq\frac{1}{n}$ and $k\geq(2+\alpha)\log(1/\varepsilon)$. There exists an explicit strong $(k,k^{\prime},\varepsilon)$-condenser
$\mathsf{RSCond}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$
@@
where $\ell=\mathopen{}\mathclose{{}\left(1+\frac{1}{\alpha}}\right)\log\frac{nk}{\varepsilon}=O_{\alpha}(\log\frac{1}{\varepsilon})$, $m=k$, and $k^{\prime}=\frac{k-\log(1/\varepsilon)}{1+\alpha}+\ell\geq(1-\alpha)k$. Note that the output entropy rate satisfies $\frac{k^{\prime}}{m}\geq 1-2\alpha$. Moreover, given $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$, $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, and a primitive element for $\mathds{F}_{2^{\ell}}$, the output $\mathsf{RSCond}(x,y)$ can be computed in time $\widetilde{O}(n)$.
In particular, if $\varepsilon^{\prime}=\sqrt{\varepsilon}$ and $k\geq\frac{\log(1/\varepsilon)}{\alpha(1+2\alpha)}$, then for all $(n,k)$-sources $X$ and a $(1-\varepsilon^{\prime})$-fraction of seeds $y$ it holds that $\mathsf{RSCond}(X,y)\approx_{\varepsilon^{\prime}}Z_{y}$, where $Z_{y}$ is an $(m=k,k^{\prime}-\ell\geq(1-2\alpha)m)$-source. Note that the seed length is $\ell=O_{\alpha}(\log\frac{1}{\varepsilon^{\prime}})$.
@@
We set $q=2^{\ell}$, and let $\zeta\in\mathds{F}_{q}$ be the generator of the multiplicative group $\mathds{F}_{q}^{\star}$ given to us as input. [^16] Working with fields of characteristic $2$ is not necessary, but may help in efficiently computing $\zeta$. For example, Shoup [Sho90] showed that given an irreducible polynomial $f\in\mathds{F}_{2}[X]$ of degree at most $d-1$, there exists a primitive element $h\in\mathds{F}_{2}[X]/\langle f\rangle$ of $\mathds{F}_{2}[X]/\langle f\rangle\equiv\mathds{F}_{2^{d}}$ such that $h$ is a monic polynomial of degree $O(\log d)$. Given $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{\ell}$, similarly to Theorem 3.10, interpret $x$ as a univariate polynomial $f_{x}$ of degree at most $d-1$, and $y$ as an element of $\mathds{F}_{q}$. The output $\mathsf{Cond}(x,y)$ is the sequence of evaluations
@@
$\mathopen{}\mathclose{{}\left(f(y),f(\zeta y),\ldots,f(\zeta^{m^{\prime}}y)}\right),$
where $m^{\prime}=\frac{m}{\log q}$.
@@
The correctness proof, as well as the exact choice of parameters, are given in [GUV09, Section 6], so we proceed to bounding the runtime. Towards that end, since we rely on a specific primitive element $\zeta$, we assume that the irreducible polynomial used to construct $\mathds{F}_{q}$ is known, either (and there are several ). Computing the evaluation points $y,\zeta y,\ldots,\zeta^{m^{\prime}}y$ can then be done naively in time $m^{\prime}\cdot M_{q}^{\mathsf{b}}(\ell)=\widetilde{O}(n)$. Then, using Lemma 2.1, the evaluation can be done in time $\widetilde{O}(d)\cdot\log q=\widetilde{O}(n)$ as well.
The “In particular” part of the theorem statement follows analogously to that of Theorem 3.10, using also the fact that if $k\geq\frac{\log(1/\varepsilon)}{\alpha(1+2\alpha)}$, then $k^{\prime}-\ell=\frac{k-\log(1/\varepsilon)}{1+\alpha}\geq(1-2\alpha)k=(1-2\alpha)m$. ∎
4 A Faster Instantiation of Trevisan’s Extractor
Definition 4.1 (weak design).
$\sum_{j<i}2^{|S_{i}\cap S_{j}|}\leq\rho(m-1).$
$\mathsf{Tre}(x,y)=\bar{x}|_{y_{S_{1}}}\circ\ldots\circ\bar{x}|_{y_{S_{m}}},$ (1)
where we denote $\bar{x}=\mathcal{C}(x)$ and interpret each length-$\log\bar{n}$ bit-string $y_{S_{i}}$ as a location in $[\bar{n}]$. For the runtime analysis, it will be important to recall that $\delta$ is set to be $\frac{\varepsilon}{cm}$ for some universal constant $c$.
Theorem 4.2.
Trevisan’s extractor of Equation 1, set to extract $m$ bits with any error $\varepsilon>0$, is computable in time $\widetilde{O}(n+m\log(1/\varepsilon))$.
@@ 4 A Faster Instantiation of Trevisan’s Extractor
Looking at Equation 1, note that we only need to compute $m$ coordinates of $\mathcal{C}(x)$. To compute those $m$ coordinates, $y_{S_{1}},\ldots,y_{S_{m}}$, we first need to compute the weak design itself. Note that this can be seen as a preprocessing step, since it only depends on the parameters of the extractor, and not on $x$ or $y$. We will use the following result.
Claim 4.3 ([FYEC24], Section A.5).
Once we have our preprocessing step, we are left with computing the code. By Corollary 3.3, we can choose $\bar{n}$ so that $n/\bar{n}=\delta^{c}$ for some universal constant $c$, and so $\bar{n}=n\cdot\operatorname{poly}(m,1/\varepsilon)$ and $\ell=\log\bar{n}=O(\log(n/\varepsilon))$. Generating the design can then be done in time $\widetilde{O}(m\log(n/\varepsilon))$. Now, Corollary 3.3 tells us that any $m$ bits of $\mathcal{C}(x)$ can be computed in time
$\widetilde{O}(n)+O(m\log(1/\delta)\log n\operatorname{loglog}n)=\widetilde{O}(n+m\log(1/\varepsilon)).$
On a RAM in the logarithmic cost model, we can use the variant of $\mathcal{C}$ that uses Spielman’s code as a base code (see Remark 3.4) and get a runtime of $O(n)+m\log(1/\varepsilon)\cdot\operatorname{polylog}(n)$. This gives a truly linear time construction whenever $m$ is at most $\frac{n}{\log(1/\varepsilon)\operatorname{polylog}(n)}$. ∎
Corollary 4.4.

### v2

@@ 3.3.2 The Lossy RS Condenser
The downside of theorem˜3.14 is that it requires the entropy in the source to be $\Omega(\log^{2}(n/\varepsilon))$, instead of the optimal $\Omega(\log(n/\varepsilon))$. Instead, we can use a lossy condenser [^21] Our extractor will lose a small constant fraction of the entropy, so losing a small constant fraction of the entropy in the condensing step will not make much difference. based on Reed–Solomon codes. Unfortunately, this comes at the expense of computing a primitive element of a field of size $\operatorname{poly}(n/\varepsilon)$, which we do not know how to do in nearly-linear time in $n$ for arbitrary $\varepsilon$-s. As in section˜3.3.1, we consider it a one-time preprocessing step, as it does not depend on the inputs to the condenser. Luckily, we have freedom in choosing the field size, as long as it is large enough. Therefore, if we choose the field size to be prime (as opposed to a power of $2$ as in [GUV09]), then we can implement this one-time preprocessing step in time $\operatorname{polylog}(n/\varepsilon)$ using randomness, as discussed in remark˜3.16.
We first state the lossy condenser over a prime field $\mathds{F}_{q}$ and argue about its running time. This condenser can be converted to a condenser over bits with only a negligible loss in parameters as in section˜3.3.1.
Theorem 3.15 (the lossy RS condenser over $\mathds{F}_{q}$ [GUV09, adapted]).
@@
For any constant $\alpha\in(0,1)$ there exist constants $C_{\alpha},C^{\prime}_{\alpha}>0$ such that the following holds for every $n\in\mathds{N}$, $\varepsilon>0$, $k\geq C_{\alpha}\log\mathopen{}\mathclose{{\left(n/\varepsilon}}\right)$, and $q$ a prime satisfying $\frac{C^{\prime}_{\alpha}}{2}\log(n/\varepsilon)\leq\log q\leq C^{\prime}_{\alpha}\log(n/\varepsilon)$. There exists a strong $(k,k^{\prime},\varepsilon)$-condenser
$\mathsf{RSCond}^{\prime}\colon\mathds{F}_{q}^{n^{\prime}}\times\mathds{F}_{q}\rightarrow\mathds{F}_{q}^{m^{\prime}}$
@@
where $n^{\prime}=\mathopen{}\mathclose{{\left\lceil\frac{n}{\lfloor\log q\rfloor}}}\right\rceil$, $m^{\prime}=\lceil k/\log q\rceil$, and $k^{\prime}\geq(1-\alpha)k$. Moreover, given $x\in\mathds{F}_{q}^{n^{\prime}}$, $y\in\mathds{F}_{q}$, and a primitive element $\zeta$ for $\mathds{F}_{q}$, the output $\mathsf{RSCond}^{\prime}(x,y)$ can be computed in time $\widetilde{O}(n)$.
@@
Given $x\in\mathds{F}_{q}^{n^{\prime}}$ and $y\in\mathds{F}_{q}$, similarly to theorem˜3.14, interpret $x$ as a univariate polynomial $f\in\mathds{F}_{q}[X]$ of degree at most $n^{\prime}-1$. Let $\zeta$ be the primitive element of $\mathds{F}_{q}$ given to us. The output $\mathsf{RSCond}^{\prime}(x,y)$ is the sequence of evaluations
@@
$\mathopen{}\mathclose{{\left(f(y),f(\zeta y),\ldots,f(\zeta^{m^{\prime}-1}y)}}\right)\in\mathds{F}_{q}^{m^{\prime}}.$
@@
The correctness proof, as well as the exact choice of parameters, are given in [GUV09, Section 6.1]. We focus on bounding the runtime. Computing the evaluation points $y,\zeta y,\ldots,\zeta^{m^{\prime}-1}y$ can be done naively in time $m^{\prime}\cdot M_{q}^{\mathsf{b}}(1)=\widetilde{O}(n^{\prime}\log q)=\widetilde{O}(n)$. Then, using lemma˜2.1, the evaluation can be done in time $\widetilde{O}(n^{\prime})\cdot\log q=\widetilde{O}(n^{\prime}\log q)=\widetilde{O}(n)$ as well. ∎
Remark 3.16 (complexity of the RS condenser preprocessing).
We now discuss the complexity of the preprocessing step required by the RS condenser, which corresponds to finding a primitive element of a field $\mathds{F}_{q}$ for a prime $q\leq\operatorname{poly}(n/\varepsilon)$. We do not know any algorithms for finding primitive elements of $\mathds{F}_{q}$ running in time $\operatorname{polylog}(q)$, unless we have access to the prime factorization of q1q-1.222222If we have access to the prime factorization of q1q-1, then we can find a primitive element of $\mathds{F}_{q}$ in time $\operatorname{polylog}(q)$ using randomness by repeatedly sampling a uniformly random element $\alpha$ of $\mathds{F}_{q}$ and checking whether it is primitive by seeing whether $\alpha^{\frac{q-1}{p}}\neq 1$ for every prime factor $p$ of q1q-1. An alternative algorithm is analyzed in [SHO05, Section 11.1]. However, the RS condenser from [GUV09] can be instantiated with any field $\mathds{F}_{q}$ of appropriately large order $q=\operatorname{poly}(n/\varepsilon)$. We can leverage this to speed up the preprocessing considerably by moving away from the fields of characteristic $2$ used in [GUV09]. More precisely, we can focus on prime $q$ and follow an approach used in cryptography for generating public parameters (prime/primitive element pairs $(q,g)$) for cryptographic schemes based on discrete logarithms over $\mathbb{Z}^{*}_{q}$, outlined in Shoup’s excellent book [SHO05, Section 11.1]. This leads to a randomized algorithm running in time $\operatorname{polylog}(n/\varepsilon)$ and succeeding with high probability, that only needs to be executed once for a given set of extractor parameters $(n,m,k,\varepsilon)$. More precisely, we can first efficiently sample the prime factorization of a uniformly random number $r$ in the set $\{1,\dots,L\}$ for an appropriate upper bound $L=\operatorname{poly}(n/\varepsilon)$, following approaches of Bach [BAC88] or Kalai [KAL03] (as discussed in [SHO05, Section 9.6]). Then, we check whether $r\geq L/2$ and $q=r+1$ is prime, and repeat if not. Since roughly a $\frac{1}{\operatorname{polylog}(n/\varepsilon)}$ fraction of such $q$’s is prime, we will succeed with high probability after $\operatorname{polylog}(n/\varepsilon)$ trials. After the process above we hold a suitably large prime $q$ and the prime factorization of q1q-1, which allows us to find a primitive element of $\mathds{F}_{q}$ in randomized time $\operatorname{polylog}(q)=\operatorname{polylog}(n/\varepsilon)$.
Finally, combining theorem˜3.15 and lemma˜3.12 allows us to get a version of the “prime qq” RS condenser over bits.
Theorem 3.17 (the lossy RS condenser over bits, [GUV09]).
For any constant $\alpha\in(0,1)$ there exist constants $C_{\alpha},C^{\prime}_{\alpha}>0$ such that the following holds for every $n\in\mathds{N}$, $\varepsilon>0$, and $k\geq C_{\alpha}\log(n/\varepsilon)$. There exists a strong $(k,k^{\prime},\varepsilon)$-condenser
$\mathsf{RSCond}\colon\{0,1\}^{n}\times\{0,1\}^{\ell}\to\{0,1\}^{m}$
where $\ell\leq C^{\prime}_{\alpha}\log(n/\varepsilon)$, $k\leq m\leq(1+\alpha)k$, and $k^{\prime}\geq(1-\alpha)m$. Moreover, given $x\in\{0,1\}^{n}$, $y\in\{0,1\}^{\ell}$, and a primitive element $\zeta$ for $\mathds{F}_{q}$ with $q$ a prime in $2^{\ell},2^{\ell+1}]$, the output $\mathsf{RSCond}(x,y)$ can be computed in time $\widetilde{O}(n)$.
In particular, if $\varepsilon^{\prime}=\sqrt{\varepsilon}$ and $C_{\alpha}>0$ is chosen large enough compared to $C^{\prime}_{\alpha}$, then for all $(n,k)$-sources $X$ and a $(1-\varepsilon^{\prime})$-fraction of seeds $y$ it holds that $\mathsf{RSCond}(X,y)\approx_{\varepsilon^{\prime}}Z_{y}$, where $Z_{y}$ is an $(m,k^{\prime}-\ell\geq(1-2\alpha)m)$-source.
@@
We argue about the output length and the output min-entropy rate. The rest is analogous to the proof of theorem˜3.14. Fix a target $\alpha\in(0,1)$. We invoke theorem˜3.15 with an appropriately small constant $\beta\in(0,1)$ in place of $\alpha$. Regarding the output length $m$, note that $m=m^{\prime}\cdot\lceil\log q\rceil\geq\frac{k}{\log q}\cdot\log q=k$, and that
$m=m^{\prime}\lceil\log q\rceil\leq m^{\prime}\log q+m^{\prime}\leq k+\log q+m^{\prime}\leq k+(\ell+1)+(k/\ell+1)=\mathopen{}\mathclose{{\left(1+\frac{1}{\ell}+\frac{\ell+2}{k}}}\right)k.$
The quantity on the right hand side can be made at most $(1+\beta)k\leq(1+\alpha)k$ by setting $\beta<\alpha$ and $C_{\beta}$ to be sufficiently larger than $C^{\prime}_{\beta}$, and by the lower bound on $\ell=\lfloor\log q\rfloor$ in theorem˜3.15. Regarding the output entropy $k^{\prime}$, since lemma˜3.12 implies a penalty of $\ell$ bits in the input min-entropy, from theorem˜3.15 instantiated with $\beta$ and lemma˜3.12 we get the guarantee that
$k^{\prime}\geq(1-\beta)(k-\ell)\geq(1-\beta)k-\ell\geq(1-2\beta)k\geq\frac{(1-2\beta)m}{1+\beta}\geq(1-\alpha)m$
provided that $C_{\beta}$ is sufficiently larger than $C^{\prime}_{\beta}$ and that $\beta\leq\alpha/3$.
The “In particular” part of the theorem statement follows analogously to that of theorem˜3.14, since $m\geq k$ and we can assume that $\ell/k\leq\alpha$ by setting $C_{\alpha}$ to be sufficiently large compared to $C^{\prime}_{\alpha}$. ∎
3.3.3 Towards removing preprocessing?
We discuss an approach, suggested by Jesse Goodman, towards removing the preprocessing steps required for the KT and RS condensers, or at least expanding the range of parameters for which preprocessing is not required, and some barriers we face when trying to fully implement this strategy.
KT condenser.
As discussed in remark˜3.11, the KT condenser in theorem˜3.14 requires a one-time preprocessing step independent of the source and the seed whenever, roughly, $\varepsilon\leq 2^{-Cn^{0.12}}$. If this holds, a sufficient preprocessing consists of generating a prime $q=\operatorname{poly}(n/\varepsilon)$, which can be done in randomized time $\operatorname{polylog}(n/\varepsilon)$. This raises the following possibility: perhaps we can derandomize this algorithm so that it only uses $O(\log(n/\varepsilon))$ bits of randomness without hurting the running time too much. If this is the case, then the randomness used by the algorithm can be absorbed into the seed, and we get a condenser running in time $\widetilde{O}(n+\operatorname{polylog}(n/\varepsilon))$, without preprocessing. When $\varepsilon$ is not too small (depending on the exponent of the $\operatorname{polylog}(n/\varepsilon)$ term), this is $\widetilde{O}(n)$.
We can realize this approach by combining the randomized prime generation algorithm with an averaging sampler. A similar strategy was used in [BIW06]. We sketch how this can be done. Let $N=\operatorname{poly}(n/\varepsilon)$. The randomized prime generation algorithm independently samples $O(\log N)$ uniformly random integers $q\in[N/2,N]$, and checks whether at least one of these numbers is prime. Primality can be tested in deterministic time $\widetilde{O}(\log^{6}N)$ using an improved variant of the AKS primality test [LEN02], and, by the prime number theorem, with high enough probability there will be at least one prime number among the samples. One way to derandomize this algorithm is by using a $(\gamma,\theta)$-averaging sampler $\mathsf{Samp}\colon\{0,1\}^{r}\to[N]^{m}$ with accuracy $\theta=\Theta(1/\log N)=\Theta(1/\log(n/\varepsilon))$ and confidence $\gamma=\varepsilon$ to generate $m$ candidate primes, and then run the AKS primality test on each candidate. Denoting by $T$ the runtime of $\mathsf{Samp}$, this procedure runs in time $\widetilde{O}(T+m\cdot\log^{6}(n/\varepsilon))$, uses $r$ bits of randomness, and generates the desired prime except with probability at most $\varepsilon$.
With some hindsight, it turns out that this running time is not enough to beat the runtime of the deterministic preprocessing from remark˜3.11, even ignoring $T$ and using a sampler with optimal sample complexity $m$. However, it is possible to improve on the runtime by replacing the AKS primality test with a faster probabilistic test, such as Miller-Rabin [SHO05, Section 10.2], and use a second sampling round to derandomize these probabilistic tests. More precisely, one iteration of the Rabin-Miller algorithm for testing primality of an integer $q$ samples $\alpha$ uniformly at random from $[q]$ and checks whether $\alpha$ passes a certain test in time $\widetilde{O}(\log^{2}q)$. [^23] Shoup [SHO05, Section 10.2] states that the test is computable in time $\widetilde{O}(\log^{3}q)$ via repeated squaring, but more sophisticated techniques yield the $\widetilde{O}(\log^{2}q)$ bound [NAR14]. When $q$ is prime all $\alpha$-s pass the test, while for $q$ composite $\alpha$ fails the test with probability at least $3/4$. We use a $(\gamma^{\prime},\theta^{\prime})$-averaging sampler $\mathsf{Samp}^{\prime}\colon\{0,1\}^{r^{\prime}}\to[N]^{m^{\prime}}$ with $\theta^{\prime}=1/4$ and $\gamma^{\prime}=\varepsilon/m$ to generate $\alpha_{1},\dots,\alpha_{m^{\prime}}$, which will be used to run Miller-Rabin tests on the prime candidates $q_{1},\dots,q_{m}$ generated by the first sampler. [^24] Note that $\mathsf{Samp}^{\prime}$ outputs samples from $[N]$, while the guarantees for the Miller-Rabin test on $q$ hold for $\alpha$ uniformly random over $[q]$. But since each $q_{i}\in[N/2,N]$, taking the test to be $\alpha\ \mathrm{mod}\ q_{i}$ for $\alpha$ uniformly random over $[N]$ only decreases the failure probability from at least $3/4$ to at least $1/2$. Note that for any fixed composite $q_{i}$, the probability that all Miller-Rabin tests $\alpha_{1}\ \mathrm{mod}\ q_{i},\dots,\alpha_{m^{\prime}}\ \mathrm{mod}\ q_{i}$ pass is at most $\varepsilon/m$. By a union bound over $q_{1},\dots,q_{m}$, the probability that this holds for at least one such $q_{i}$ is at most $m\cdot\varepsilon/m=\varepsilon$. Therefore, if $\mathsf{Samp}$ and $\mathsf{Samp}^{\prime}$ run in time $T$ and $T^{\prime}$, respectively, this procedure runs in time $\widetilde{O}(T+T^{\prime}+m\cdot m^{\prime}\cdot\log^{2}(n/\varepsilon))$, uses $r+r^{\prime}$ bits of randomness, and generates the desired prime except with probability at most $2\varepsilon$.
We already know from remark˜3.11 that the KT condenser does not require preprocessing to run in $\widetilde{O}(n)$ time when $\varepsilon>2^{-Cn^{0.12}}$, so it is relevant to explicitly work out the best possible improvement afforded by the approach above, assuming we are aiming for $\widetilde{O}(n)$ runtime. We can instantiate $\mathsf{Samp}$ and $\mathsf{Samp}^{\prime}$ with the nearly-optimal efficient averaging samplers of Xun and Zuckerman [XZ25, Theorem 2]. Under the choices of $N$, $(\theta,\gamma)$, and $(\theta^{\prime},\gamma^{\prime})$ above, for any constant $\delta>0$, we can instantiate $\mathsf{Samp}$ with randomness complexity $r=O_{\delta}(\log(n/\varepsilon))$ and sample complexity $m=O(\log^{3+\delta}(n/\varepsilon))$, and $\mathsf{Samp}^{\prime}$ with randomness complexity $r^{\prime}=O_{\delta}(\log(n/\varepsilon))$ and sample complexity $m^{\prime}=O(\log^{1+\delta}(n/\varepsilon))$.
Therefore, using these samplers, for any constant $\delta>0$, the prime-generating procedure above uses $O_{\delta}(\log(n/\varepsilon))$ bits of randomness, which can be absorbed into the seed of the condenser, runs in time $\widetilde{O}(T+T^{\prime}+\log^{6}(n/\varepsilon))$, with $T$ and $T^{\prime}$ the runtimes of $\mathsf{Samp}$ and $\mathsf{Samp}^{\prime}$, and fails with probability at most $2\varepsilon$, which can be absorbed into the error of the condenser. One can verify that $T$ and $T^{\prime}$ are $O(\log^{6}(n/\varepsilon))$ in this regime. Therefore, the time bound above is $\widetilde{O}(n)$ when $\varepsilon>2^{-Cn^{\gamma}}$ for any constant $\gamma<1/6$, which improves on the simpler $\varepsilon>2^{-Cn^{0.12}}$ bound from remark˜3.11.252525In contrast, using the AKS primality test would lead to running time $\widetilde{O}(\log^{9}(n/\varepsilon))$, which is only $\widetilde{O}(n)$ when $\varepsilon>2^{-Cn^{1/9}}$, and therefore worse than the bound from remark 3.11. We also note that we cannot hope to improve on this bound by picking a better averaging sampler, since any $(\theta,\gamma)$-averaging sampler must have sample complexity $m=\Omega(\log(1/\gamma)/\theta^{2})$ [CEG95].
RS condenser.
Unlike the KT condenser, the RS condenser from theorem˜3.17 always requires preprocessing. Therefore, arguably the most interesting direction in this discussion would be to establish a statement of the form “if $\varepsilon\geq 2^{-n^{\gamma}}$ for some constant $\gamma>0$, then the RS condenser does not require preprocessing to run in time $\widetilde{O}(n)$”, in analogy with the statement we already have for the KT condenser. However, it is not clear to us how to implement a sampler-based strategy in this case. We elaborate on this below.
The preprocessing algorithm considered in remark˜3.16 has two stages: (1) repeatedly sample an integer $q$ alongside its prime factorization uniformly at random from $[N/2,N]$, with $N=\operatorname{poly}(n/\varepsilon)$, until $q+1$ is prime, as in [SHO05, Section 9.6], and (2) find a primitive element of $\mathds{F}_{q}$. We focus on the first stage. The first part of this stage requires sampling a “random non-increasing sequence” in $[1,N]$ (see [SHO05, Section 9.5 (Algorithm RS)]). However, this procedure requires too much randomness, which blows up the seed length of the sampler. Indeed, following the analysis in [SHO05, Section 9.5.2] gives that Algorithm RS requires $\Theta(\log^{2}N)=\Theta(\log^{2}(n/\varepsilon))$ bits of randomness in expectation. More precisely, let $O_{i}$ be the random variables denoting the number of times the integer $i$ appears in the non-increasing sequence, as defined in [SHO05, Section 9.5.2]. There, it is shown that $\mathds{E}[O_{i}]=\frac{1}{i-1}$ for all integers $i\in[2,N]$. Every time $i$ is sampled requires using at least $\log i$ bits of randomness in that iteration of Algorithm RS. Therefore, the expected number of bits of randomness required is at least
$\sum_{i=1}^{N}\log i\cdot\mathds{E}[O_{i}]\geq\sum_{i=1}^{N}\frac{\log i}{i}\geq\int_{1}^{N}\frac{\log x}{x}\;dx-O(1)=\Theta(\log^{2}N).$
Therefore, the sampler must output samples of bitlength $\Omega(\log^{2}(n/\varepsilon))$, and so also requires a seed of length $\Omega(\log^{2}(n/\varepsilon))$.
We note that the barriers outlined above are specific to the preprocessing steps we considered in this work, and to the use of samplers. It may be possible to improve on the current results by considering alternative preprocessing steps and by using other pseudorandom objects. We leave this as a natural direction for future work.